Suppose three test kits are tested at random. Let N be the random variable for the number of non - defective test kits, construct the probability distribution P(N). What can you conclude about the effectivity of the COVID-19 Rapid Antibody Test kits based on your probability distribution? Explain briefly.



1
Expert's answer
2022-02-22T08:38:25-0500

Solution:

Let X1, X2 and X3 are three kits that are tested random

"X_{i,\ i=1,2,3}=1\ if\ X_i=D
 \ \ \ \ \\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ =0\ if \ X_i=N"

Let, total number of test kits that are defective =d 

and total number of test kits that are non-defective =n

"\therefore P[x_ i=D]=\frac{d}{d+n}=p(\operatorname{let})"

Then, "x=\sum_{i=1}^{3} x i="  number of defective test kits.

Now, "x_{i} \sim"  bernoulli (p)

"\therefore \sum_{i=1}^{3} x i \sim"  binomial (3, p) where, "p=\frac{d}{d+n}"

"\begin{aligned}

\therefore P[X=x] &=\left(\begin{array}{l}

3 \\

x

\end{array}\right) p^{x}(1-p)^{3-x}, x=0,1,2,3 \\

&=0 \text { otherwise. }

\end{aligned}"